OpenStudy (anonymous):

find the integral S (x/6 + 6/x) dx

OpenStudy (anonymous):

You can break apart the integral

OpenStudy (anonymous):

\[\int\limits_{}^{}({x \over 6}+{6 \over x})dx={x^2 \over 12}+6\ln \left| x \right|+c={x^2 \over 12}+\ln (x^6) +c\]

OpenStudy (anonymous):

how did u get x^12/12?

OpenStudy (anonymous):

It's x^2/12. Because when you integration of x is x^2/2, then multiply this by 1/6, you get x^2/12.

OpenStudy (anonymous):

hey, anwarA. How did yu do for the fraction to appear like that ?

OpenStudy (anonymous):

How much would you pay for that? :P

OpenStudy (amistre64):

frac{}{}

OpenStudy (anonymous):

i'll givu a medal haha

OpenStudy (anonymous):

@amistre64: he/she asked me >.<

OpenStudy (amistre64):

lol....then answer ;) if you know how lol

OpenStudy (anonymous):

oh thanks amistre \[frac{1}{2}\]

OpenStudy (anonymous):

phail lol

OpenStudy (amistre64):

i left out that part where you do it in the....

OpenStudy (anonymous):

I'll tell you another way, a better way. "{something over something}"

OpenStudy (anonymous):

using the "equation" feature of course.

OpenStudy (anonymous):

\[1 \over 2\] oh boy thanks lo

OpenStudy (anonymous):

You are welcome :)